OpenStudy (anonymous):

OlWbat is the sol 2x-10y+6z= 28 -x+2-z=0 10x-y-z=21

OpenStudy (anonymous):

-x+2y-z=0?

OpenStudy (unklerhaukus):

solve the second equation for z

OpenStudy (anonymous):

I do t hVw tome

OpenStudy (anonymous):

2x-10y+6z= 28 -x+2y-z=0 10x-y-z=21 is that the question?

OpenStudy (anonymous):

@Mandy86 u there?

OpenStudy (anonymous):

-28

OpenStudy (anonymous):

2x-10y+6z= -28 -x+2y-z=0 10x-y-z=21 corretc

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

2x-10y+6z= -28 -x+2y-z=0 10x-y-z=-21 anything else

OpenStudy (anonymous):

stupid phone wouldnt work.

OpenStudy (anonymous):

is the above correct

OpenStudy (anonymous):

dont worry about it i submitted it allready

OpenStudy (anonymous):

what did u get?